The measure of the angles of a triangle are x,2x, and 3x. What are the measures of the angles of the triangle?
Answer by checkley75(3666)

THE SUM OF THE THREE ANGLES OF A TRIANGLE=180 DEGREES. THUS
X+2X+3X=180
6X=180
X=180/6
X=30 WHICH IS THE SMALLEST ANGLE. THE OTHER TWO ARE
2X=2*30=60 &
3X=3*30=90
PROOF
30+60+90=180
180=180